A market research firm reported that the mean annual earning of all family practitioners in the United States is $200,000. A public health researcher is concerned that Harrisburg family practitioners earn less than this national average. They take random sample of 55 family practitioners in the Harrisburg, PA area and found a test statistic of t = -1.24. Calculate the p-value for this hypothesis test at the alpha=0.01 level, assuming the population standard deviation is unknown and the original distribution is approximately normal. Round the p-value to 4 decimal places.
Answer #1

Here is given that test statistic t = -1.24 and n = 55

Degrees of freedom = n - 1 = 55 -1 = 54

The p - value for this two tailed hypothesis test at the alpha = 0.01 level is " 0.2203 "